OpenStudy (anonymous):

How many multiples of 3 are there between 1167 and 2048 inclusive?

OpenStudy (anonymous):

I have started off with Tn = 1167 + (n-1)3 = 1167 + 3n-3

OpenStudy (anonymous):

im stuck after that

jimthompson5910 (jim_thompson5910):

Notice how 2048 isn't a multiple of 3, but 2046 is since 2046/3 = 682 So 1167/3 = 389 and 2046/3 = 682 which means 1167 = 3*389 and 2046 = 3*682

OpenStudy (anonymous):

Your first term is 1167 and last term is 2048 with common difference 3

jimthompson5910 (jim_thompson5910):

So the sequence 1167, 1170, ..., 2046 can be written as 3*389, 3*390, ..., 3*682

jimthompson5910 (jim_thompson5910):

So you need to count how many numbers there are from 389 to 682

OpenStudy (anonymous):

293?

jimthompson5910 (jim_thompson5910):

close, 294

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

thanks